I am trying to get the correct answer to calculation of capacity for:
B/240+(B*2)=12/27
I have the answer as B = 960, but, I need the step by step solving of the problem. Thanks.

Answer by ldjackson(1)

12/27 = B/240+(B*2)
***Cross Multiply with "12/27" and the "B" being divided on the right
B = (12/27*240) + (12/27*2B)
B = 106.66666 + .88888B
.11112B = 106.66666
B = 959.92